Consider the following aggregate expenditure model of the Canadian economy operating with given wages and other factor prices, price level, interest rates, exchange rates, and expectations:
C = 50 + 0.8YD I = 400 G = 500 T = 0.3Y X = 650 IM = 0.36Y
where C is consumption (the 0.8 term represents the marginal propensity to consume) YD is disposable income, I is investment, G is government spending on goods and services, T is the total value of taxes net of transfers (the 0.3 term represents the net tax rate on national income), X is exports, and IM is imports (the 0.36 term represents the marginal propensity to import).
(b) Calculatethelevelofdisposableincome,consumption,privatesaving,governmentbudgetbalance, and net exports at the equilibrium. Express the components of aggregate expenditure (C, I, G, and NX) as percentages of GDP (to one decimal place).
1
Expert's answer
2015-02-26T09:26:10-0500
C = 50 + 0.8YD
I = 400
G = 500
T = 0.3Y
X = 650
IM = 0.36Y
(b) Y = C + I + G + NX
C = 50 + 0.8YD = 50 + (c - im)*(Y - T) = 50 + 0.44*0.7Y = 50 + 0.308Y
Y = 50 + 0.308Y + 400 + 500 + 650 - 0.36Y
1.052Y = 1600
Y = $1520.9
Disposable income YD = 0.308*1520.9/0.8 = 585.5
Consumption C = 50 + 0.308*1520.9 = 518.4
Private saving Sp = Y - T - C = 0.7*1520.9 - 518.4 = 546.2
Government budget balance BS = T - G = 0.3*1520.9 - 500 = -43.7
Net exports NX = X - IM = 650 - 0.36*1520.9 = 102.5
The components of GDP are:
C = 518.4/1520.9*100% = 34.1%
I = 400/1520.9*100% = 26.3%
G = 500/1520.9 = 32.9%
NX = 6.7%